# HeMiTo-dynamics: a characterisation of mammalian prion toxicity using non-dimensionalisation, linear stability and perturbation analyses

### v1

@@ 7 Discussion and conclusions
By constructing and analysing a class of mechanistic models of prion-like replication, we suggest that the accumulation of toxic prions in mammalian cells over time is characterised by three phases referred to as HeMiTo-dynamics. Given that the initial concentrations of prions are low, mammalian cells initially form healthy prions during the He-phase while there is no formation of the toxic form. Importantly, this phase is identical for all types of conversion functions $f$ as Theorem 3 suggests, and the main aim of this phase is to build up the levels of the healthy form. When a sufficiently high concentration of the healthy form is reached, the system enters the Mi-phase, where the healthy form reaches a maximum value while the toxic form increases exponentially. Here, the choice of conversion function $f$ determines the qualitative behaviour of the trajectories of the healthy form $u$ during the Mi-phase, and our asymptotic approximation $u_{\mathrm{Mi}}$ in Eq. (38) indicates that there are two main types of dynamical behaviour during this phase. On the one hand, trajectories in the Mi-phase can behave like the SIR model where $u_{\mathrm{Mi}}(\tau)$ is a concave function which reaches a clear maximum point $u_{\max}$ before it decreases, and the reference model by Thompson et al. [16] defined by $f(v)=1.00$ has this type of SIR-like behaviour in the Mi-phase (FIG. 4A). On the other hand, trajectories for the healthy form in the Mi-phase can also behave like a logistic growth function where the maximum value $u_{\max}$ corresponds to the carrying capacity of $u_{\mathrm{Mi}}(\tau)$, and, for instance, the model defined by $f(v)=0.57+v^{4}/(1+v^{5})$ has this type of dynamical behaviour (FIG. 4C). Comparing these two cases, the formation of toxic prions, which occurs at the expense of the healthy form, is slower for the latter logistic growth type of dynamics compared to the SIR-like counterpart. This is also clear from the approximate system of ODEs in Eqs. (28) and (29) describing the dynamics of our prion models during the Mi-phase, as the conversion rate is essentially given by $f(v_{0})$, i.e. the value of the conversion function $f$ evaluated at the initial concentration of the toxic form $v_{0}$. In our simulations where $v_{0}=0.05$, it is clear that the value of $f(v_{0})$ is much lower for the function $f(v)=0.57+v^{4}/(1+v^{5})$ compared to the reference model $f(v)=1.00$ (FIG. 1) and thus it is not surprising that the time to reach the TSS is much longer in the former case compared to the latter (FIG. 2).
@@ Appendix B Proof of Theorem 16: existence of steady states
00 0 $=\varepsilon{v^{\star}}\left(\varepsilon{u^{\star}}f(v^{\star})-1\right)\,.$ (41)
@@ Appendix C Proof of Proposition 18: condition for the HSS being a saddle point
and its trace and determinant are given by
$\mathrm{Tr}(\mathcal{J}(u,v))$ $=-(c_{2}+\varepsilon)-\varepsilon{v}f(v)+\varepsilon{u}(f(v)+vf^{\prime}(v))\,,$ (43)

### v2

@@ 7 Discussion and conclusions
By constructing and analysing a class of mechanistic models of prion-like replication, we suggest that the accumulation of toxic prions in mammalian cells over time is characterised by three phases referred to as HeMiTo-dynamics. In the healthy phase, the concentration of healthy proteins normalises to physiological levels, while there is no formation of toxic protein. Importantly, this phase is identical for all types of conversion functions $f$ as Theorem 3 suggests, and the phase is dominated by the dynamics of healthy form. When interpreting the biological meaning of the He-phase, it is important to consider that we are assuming that the model parameters are in a regime where the TSS is stable. In this specific parameter regime, the toxic form will inevitably grow to a nonzero steady-state. Suppose the initial conditions are close to the physiological levels (the peak of the healthy concentration reached at the beginning of the Mi-phase). In that case, this phase is short-lived, and we should interpret the model as beginning at a time very close to disease initiation. In other words, the Mi-phase will begin abruptly. However, the He-phase may also represent the period in time in which the system parameters have changed, altering the levels of the healthy form prior to the production of the toxic form. In a biological setting, the heterodimer model may or may not exhibit a change in physiological levels of healthy protein before the production of toxic protein dependent on which model parameter induced the pathology (lowering of toxic clearance, higher production of healthy protein, or increased in the conversion of healthy to toxic form). As physiological levels of healthy protein remain mostly unchanged, future research should focus on mechanisms of disease initiation that do not alter healthy form concentration as predicted by the heterodimer model.
@@
When a sufficiently high concentration of the healthy form is reached, the system enters the Mi-phase, where the healthy form reaches a maximum value while the toxic form increases exponentially. Here, the choice of conversion function $f$ determines the qualitative behaviour of the trajectories of the healthy form $u$ during the Mi-phase, and our asymptotic approximation $u_{\mathrm{Mi}}$ in Eq. (38) indicates that there are two main types of dynamical behaviour during this phase. On the one hand, trajectories in the Mi-phase can behave like the SIR model where $u_{\mathrm{Mi}}(\tau)$ is a concave function which reaches a clear maximum point $u_{\max}$ before it decreases, and the reference model by Thompson et al. [18] defined by $f(v)=1.00$ has this type of SIR-like behaviour in the Mi-phase (FIG. 4A). On the other hand, trajectories for the healthy form in the Mi-phase can also behave like a logistic growth function where the maximum value $u_{\max}$ corresponds to the carrying capacity of $u_{\mathrm{Mi}}(\tau)$, and, for instance, the model defined by $f(v)=0.57+v^{4}/(1+v^{5})$ has this type of dynamical behaviour (FIG. 4C). Comparing these two cases, the formation of toxic prions, which occurs at the expense of the healthy form, is slower for the latter logistic growth type of dynamics compared to the SIR-like counterpart. This is also clear from the approximate system of ODEs in Eqs. (28) and (29) describing the dynamics of our prion models during the Mi-phase, as the conversion rate is essentially given by $f(v_{0})$, i.e. the value of the conversion function $f$ evaluated at the initial concentration of the toxic form $v_{0}$, which arises from the assumption that $f(v)\sim\mathcal{O}(v)$. In our simulations where $v_{0}=0.05$, it is clear that the value of $f(v_{0})$ is much lower for the function $f(v)=0.57+v^{4}/(1+v^{5})$ compared to the reference model $f(v)=1.00$ (FIG. 1) and thus it is not surprising that the time to reach the TSS is much longer in the former case compared to the latter (FIG. 2).
@@
Comment on reproducibility
To run the analyses presented in this work, a script is available at the public github-repositry associated with this project; https://github.com/JohannesBorgqvist/HeMiTo.
@@ Appendix B Proof of Theorem 16: existence of steady states
00 0 $=\varepsilon{v^{\star}}\left({u^{\star}}f(v^{\star})-1\right)\,.$ (41)
@@ Appendix C Proof of Proposition 18: condition for the HSS being a saddle point
and its determinant by